OpenStudy (anonymous):

How do you simplify the cubed root of (25/16)?

OpenStudy (anonymous):

You wouldn't. Are you sure that isn't the squared root of? \[\sqrt{\frac{ 25 }{ 16 }}\] ?

OpenStudy (anonymous):

I wish it was... D: I get \[\sqrt[3]{25}/2\sqrt[3]{2}\] but it's not one of the answer choices and I don't know what to do from there.

OpenStudy (anonymous):

What are the other choices?

OpenStudy (anonymous):

\[\sqrt[3]{25}/4\] \[4/\sqrt[3]{100}\] \[\sqrt[3]{100}/4\] \[5/4\]

OpenStudy (anonymous):

Okay. I get what they did. You have: \[\sqrt[3]{\frac{ 25 }{ 16 }}\] If you take that fraction and multiply the top and bottom by 4 you receive: \[\sqrt[3]{\frac{ 100 }{ 64 }} \] Now then, the cubed root of 64 is 4. So you get 100^(1/3) over 4 or choice C.

OpenStudy (anonymous):

Oh I see now~ Thank you so much :)

OpenStudy (anonymous):

It isn't really simplification. It is "can you play around with numbers and get something equivalent."